OpenStudy (anonymous):

Solve this (1/4 - 7/8) / (-2 3/16). Explain how to work it

zepdrix (zepdrix):

im confused by the way this is written, is the bottom a mixed number? or is it -23/16? :o

OpenStudy (anonymous):

No its a mixed number

OpenStudy (anonymous):

-2 & 3/16

zepdrix (zepdrix):

ok ok first deal with the mixed number, noone likes those things.. get rid of it!! :U

OpenStudy (anonymous):

Ok i got -29/16

OpenStudy (anonymous):

Is that correct?

zepdrix (zepdrix):

sorry im alil rusty on mixed numbers. had to look it up a quick sec. When you see something like -2&3/16, the negative is being applied to the entire quantity. so it's -2 and -3/16.

zepdrix (zepdrix):

so -2 is -32/16 with another -3/16 would be -35/16, yes? =D

OpenStudy (anonymous):

Yes i got that

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

did you simpify the top yet? c:

OpenStudy (anonymous):

I got -5/8, not sure if that's correct

hero (hero):

I'm assuming it's this based on what you posted. \[\frac{\left(\frac{1}{4} - \frac{7}{8}\right)}{-2\frac{3}{16}}\]

OpenStudy (anonymous):

Yes that's the sum

hero (hero):

Okay, well rewrite it like this: \[\frac{1}{4} - \frac{7}{8} \div -2\frac{3}{16}\]

zepdrix (zepdrix):

yes -5/8 on top :) sounds good

hero (hero):

\[-\frac{5}{8} \div -2\frac{3}{16}\]

hero (hero):

Now what do we do?

OpenStudy (anonymous):

Change the mixed fraction to an improper fraction

OpenStudy (anonymous):

We got -35/16

hero (hero):

\[-\frac{5}{8} \div -\frac{35}{16}\] Now what?

hero (hero):

Would you like a hint?

OpenStudy (anonymous):

Yes plz

hero (hero):

Does the word reciprocal ring a bell?

OpenStudy (anonymous):

Im thinking we have to do the reciprocal & multiply

OpenStudy (anonymous):

Lol yea

hero (hero):

\[-\frac{5}{8} \times -\frac{16}{35}\] Now what do we do?

OpenStudy (anonymous):

Im getting 2/7

hero (hero):

That was easy, wasn't it?

OpenStudy (anonymous):

So thats it?

hero (hero):

You can divide 2/7, but then you would no longer have an exact value. We started with fractions, so we should end with a fraction. Once you have reduced a fraction problem to its simplest form, you're pretty much done. So yeah, "That's it". :)

OpenStudy (anonymous):

Yay!!! Thanks guys :-D

hero (hero):

You didn't admit that it was easy

OpenStudy (anonymous):

Yes it was

hero (hero):

I know it was. You did all the work.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Yes it was

OpenStudy (anonymous):

Yay!!! Thanks guys :-D

OpenStudy (anonymous):

So thats it?

OpenStudy (anonymous):

Im getting 2/7